OpenStudy (angel_kitty12):

Given f(x)=4x^3+2x^2+x-7 f(-1/2) Please help

OpenStudy (campbell_st):

just substitute x = -1/2 into the polynomial and calculate and answer... \[P(\frac{-1}{2}) = 4 \times (\frac{-1}{2})^3 + 2 \times (\frac{-1}{2})^2 + (\frac{-1}{2}) - 7\]

OpenStudy (angel_kitty12):

I tried doing that but I mixed up my calculations along the way.

OpenStudy (angel_kitty12):

I think I multiplied wrong.

OpenStudy (angel_kitty12):

I'm not very good with fractions.

OpenStudy (campbell_st):

well just type it into the calculator the way I've written it using brackets... I think the answer will be negative.

OpenStudy (angel_kitty12):

Oh okay I'll try. Thanks.